# Free Descriptions of Convex Sets

### v2

@@ 4.7 Graphon Parameters
Figure 2: Weighted undirected graph represented as a graph, an adjacency matrix $X$, and a symmetric function (graphon) $W_{X}$ on $[0,1]^{2}$.
@@
$\ +L_{4}\otimes\left(\mathrm{diag}(X)\mathbbm{1}^{\top}+\mathbbm{1}\mathrm{diag}(X)^{\top}\right)+L_{5}\otimes X+L_{6}\otimes\mathbbm{1}\mathbbm{1}^{\top}+L_{7}\otimes(2^{n}I_{2^{n}})\succeq 0\Bigg{\}},\quad L_{7}\succeq 0.$
Note that all the functions of $X$ appearing in the above description only depend on the associated step graphon $W_{X}$. For example, $\frac{\mathbbm{1}^{\top}X\mathbbm{1}}{2^{2n}}=\int_{[0,1]^{2}}W_{X}(t,s)\,\mathrm{d}t\,\mathrm{d}s$ and $\frac{\mathrm{Tr}(X)}{2^{n}}=\int_{0}^{1}W_{X}(t,t)\,\mathrm{d}t$.
Expressing the above free description in terms of graphons yields a description of a limiting convex set by Theorem 3.6. Its gauge function is, in turn, a convex graphon parameter extending continuously to the corresponding limit. Endow $\mathbb{V}_{\infty}$ with the $L^{\infty}([0,1]^{2})$-norm, noting that $\|\mathcal{P}_{n}W\|_{\infty}\leq\|W\|_{\infty}$ for all $W\in L^{\infty}([0,1]^{2})$ by Jensen’s inequality. We view elements $[W_{i,j}(x,y)]_{i,j=1}^{k}\in\mathbb{S}^{k}\otimes L^{\infty}([0,1]^{2})$ as symmetric matrices with entries in $L^{\infty}([0,1]^{2})$ (so $W_{j,i}=W_{i,j}$), and denote
$[W_{i,j}(x,y)]_{i,j=1}^{k}\succeq 0\quad\textrm{ if }\quad\int_{[0,1]^{2}}\sum_{i,j=1}^{k}W_{i,j}(x,y)f_{i}(x)f_{j}(y)\,dx\,dy\geq 0\quad\textrm{ for all }\quad f_{1},\ldots,f_{k}\in L^{2}([0,1]).$ (21)
Remarkably, convex graphon parameters that extend continuously to this limit are also projection-compatible by [79, Thm. 3.17].
@@
where $\overline{\mathbb{V}_{\infty}}\subseteq L^{\infty}([0,1]^{2})$ is a subspace of symmetric bounded measurable functions on $[0,1]^{2}$.
@@
Endow $\mathbb{U}_{\infty}=\mathbb{S}^{k}\otimes\mathbb{V}_{\infty}$ with the norm $\|[W_{i,j}]_{i,j=1}^{k}\|=\max_{i,j}\|W_{i,j}\|_{\infty}$, so $\overline{\mathbb{U}_{\infty}}=\mathbb{S}^{k}\otimes\overline{\mathbb{V}_{\infty}}$. Then $\overline{\mathcal{K}_{\infty}}$ is the cone of PSD matrices $[W_{i,j}(x,y)]_{i,j=1}^{k}\succeq 0$ in $\overline{\mathbb{U}_{\infty}}$. Furthermore, for any $W\in\mathbb{V}_{\infty}$ we have
@@ 5.1 The PSD Cone and Variants
$\mathbb{U}_{n}^{\mathsf{G}_{n}}=\left\{Q_{n}\bigoplus_{\lambda\in\Lambda}(X_{\lambda}\otimes I_{\dim\mathbb{W}_{\lambda[n]}})Q_{n}^{*}:X_{\lambda}\in\mathbb{S}^{m_{\lambda}}\right\},$ (24)
@@
$\mathrm{Sym}^{2}_{+}(\mathbb{V}_{n})^{\mathsf{G}_{n}}=\left\{X\in\mathbb{U}_{n}^{\mathsf{G}_{n}}:X\succeq 0\right\}=\left\{Q_{n}\bigoplus_{\lambda\in\Lambda_{n}}(X_{\lambda}\otimes I_{\dim\mathbb{W}_{\lambda}})Q_{n}^{*}:X_{\lambda}\in\mathbb{S}^{m_{\lambda}}_{+}\right\}.$ (25)
@@
We now prove Theorem 1.12 giving constant-sized descriptions for invariant sums-of-squares.
@@
If $v(x)$ is the vector whose coordinates are all the monomials in the $\binom{n}{k}$ variables $x_{i_{1},\ldots,i_{k}}$ of degree at most $d$, then [64, Thm. 3.39] yields
$\mathrm{SOS}_{\mathbb{U}_{n}}=\pi_{n}\left(\mathrm{Sym}_{+}^{2}\left(\mathrm{Sym}^{\leq d}\left(\bigwedge\nolimits^{k}\mathbb{R}^{n}\right)\right)\right),\quad\pi_{n}(M)=v(x)^{\top}Mv(x)+\mathcal{I}_{n}.$
The map $\pi_{n}\colon\mathrm{Sym}^{2}\left(\mathrm{Sym}^{\leq d}\left(\bigwedge^{k}\mathbb{R}^{n}\right)\right)\to\mathbb{U}_{n}$ is equivariant by definition of the action of $\mathsf{G}_{n}$ and the invariance of $\mathcal{I}_{n}$. If $\mathscr{V}_{0}=\{\mathbb{R}^{n}\}$ as in Example 1.5, then $\mathscr{V}=\mathrm{Sym}^{\leq d}\left(\bigwedge^{k}\mathscr{V}_{0}\right)$ is a $\mathscr{V}_{0}$-module generated in degree $kd$, and presented in degree $kd$ if $\mathsf{G}_{n}=\mathsf{S}_{n},\mathsf{B}_{n}$ and in degree $kd+1$ if $\mathsf{G}_{n}=\mathsf{D}_{n}$ by Theorem 2.11. Thus, the result follows from Theorem 5.1 and Proposition 5.3. ∎
@@ 5.2 The Relative Entropy Cone and Variants
Set $d_{0}=0$ if $\mathsf{G}_{n}=\mathsf{S}_{n},\mathsf{B}_{n}$ and $d_{0}=1$ if $\mathsf{G}_{n}=\mathsf{D}_{n}$. Let ^$\hat{\mathcal{B}}\subseteq\mathcal{B}_{d_{W}}$ be a set of minimal-degree $\mathsf{G}_{d_{W}+d_{0}}$-orbit representatives in $\mathcal{B}_{d_{W}+d_{0}}$, which are also orbit representatives for $\mathcal{B}_{n}$ for all $n\geq d_{W}+d_{0}$ by Proposition 2.15(c). Any $M\in\widetilde{\mathbb{U}_{n}}^{\mathsf{G}_{n}}=\mathcal{L}(\mathbb{R}^{\mathcal{B}_{n}},\mathbb{U}_{n})^{\mathsf{G}_{n}}$ for $n\geq d_{W}+d_{0}$ is fully determined by the images $Me_{\beta}\in\mathbb{U}_{n}^{\mathsf{Stab}_{\mathsf{G}_{n}}(\beta)}$ of the basis elements $e_{\beta}$ for $\beta\in\hat{\mathcal{B}}$ and conversely, for any collection $\left\{u_{\beta}\in\mathbb{U}_{n}^{\mathsf{Stab}_{\mathsf{G}_{n}}(\beta)}\right\}_{\beta\in\hat{\mathcal{B}}}$ there is a unique $M\in\widetilde{\mathbb{U}_{n}}^{\mathsf{G}_{n}}$ satisfying $Me_{\beta}=u_{\beta}$. Moreover, $M\in\widetilde{\mathcal{K}_{n}}$ if and only if $Me_{\beta}\in\mathcal{K}_{n}^{\mathsf{Stab}_{\mathsf{G}_{n}}(\beta)}$ for all $\beta\in\hat{\mathcal{B}}$. Thus, we have
@@
As an application of Proposition 5.5, we obtain constant-sized descriptions for SAGE cones of signomials. Indeed, if $\mathcal{A}_{n},\mathcal{B}_{n}\subseteq\mathbb{R}^{n}$ as in the above proposition, define the sequence =$\mathscr{F}=\{\mathbb{F}_{n}\}$ of functions on $\{\mathbb{R}^{n}\}$
@@
$\mathrm{SAGE}_{\mathcal{A}_{n},\mathcal{B}_{n}}$ =$=\Big{\{}(c,t)\in\mathbb{R}^{\mathcal{A}_{n}}\oplus\mathbb{R}^{\mathcal{B}_{n}}:\exists M=(\nu^{(\beta)},c^{(\beta)},t_{\beta})_{\beta\in\mathcal{B}_{n}}\in\mathrm{REM}_{\mathcal{A}_{n},\mathcal{B}_{n}}\textrm{ s.t.}\sum\nolimits_{\beta\in\mathcal{B}_{n}}c^{(\beta)}=c,$
$\qquad\sum\nolimits_{\alpha\in\mathcal{A}_{n}}\nu^{(\beta)}_{\alpha}(\alpha-\beta)=0\textrm{ for all }\beta\in\mathcal{B}_{n}\Big{\}}$
@@
$\mathbb{L}_{n}$ =$=\left\{M=(\nu^{(\beta)},c^{(\beta)},t_{\beta})_{\beta\in\mathcal{B}_{n}}\in\widetilde{\mathbb{U}_{n}}:\sum\nolimits_{\alpha\in\mathcal{A}_{n}}\nu^{(\beta)}_{\alpha}(\alpha-\beta)=0\textrm{ for all }\beta\in\mathcal{B}_{n}\right\},$
@@
6 Free Convex Regression
In this section, we use our framework to describe a solution to the free convex regression problem. Recall that in this problem we are given evaluation data in different dimensions and we seek a sequence of convex functions that best fit the data and can be instantiated in any desired dimension (including those not represented in the data). To that end, we use the framework we developed in Section 3 to obtain parametric families of freely-described convex sets in a fully algorithmic manner. We then present a numerical procedure to fit elements of these families to the given data. Our implementation of the resulting algorithms can be found at https://github.com/eitangl/anyDimCvxSets.
@@ 6.1 Computationally Parametrizing Descriptions
1:Input: Consistent sequences $\mathscr{V},\mathscr{W},\mathscr{U}$.
@@
Fix $n_{0}\in\mathbb{N}$ as in Algorithm 1. We now explain how to find bases for the desired spaces of equivariant linear maps $A_{n_{0}}$ and $B_{n_{0}}$ in a fixed dimension, and how to extend them to bases of freely-described elements of $\mathscr{V}\otimes\mathscr{U}$ and $\mathscr{W}\otimes\mathscr{U}$. These elaborate on steps 5, 8 and 10 in Algorithm 1.
@@
Let $\mathsf{G}$ be a real Lie group with finitely-many connected components acting on a vector space $\mathbb{V}$ via the homomorphism $\rho\colon\mathsf{G}\to\mathrm{GL}(\mathbb{V})$. Let $\{H_{i}\}$ be a basis for the Lie algebra $\mathfrak{g}$ of $\mathsf{G}$ and $\{h_{j}\}$ be a finite collection of discrete generators. Then
@@
The equations (28) express the space $\mathcal{L}(\mathbb{V}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ as the kernel of a matrix, which is often very large and sparse. A basis for the kernel of such a matrix can be computed using its LU decomposition as in [84, 85], or using the algorithm of [31, §5]. A basis for the space $\mathcal{L}(\mathbb{W}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ is obtained analogously.
@@
hence the corresponding subspace of $\mathcal{L}(\mathbb{V}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ is the kernel of the matrix obtained by combining (28)-(30). The subspace of $\mathcal{L}(\mathbb{W}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ satisfying the hypotheses of Theorem 3.5 is again the kernel of a matrix and its basis is computed similarly.
@@
Given bases of equivariant $A_{n_{0}}^{(i)},B_{n_{0}}^{(j)}$, we wish to extend them to freely-described elements. We do so computationally by applying a linear map for $n<n_{0}$ and solving a linear system for each $n>n_{0}$ to which we wish to extend. For each $n<n_{0}$, we set $A_{n}^{(i)}=\psi_{n_{0},n}^{*}A_{n_{0}}^{(i)}\varphi_{n_{0},n}$ and similarly for $B_{n}^{(j)}$. For each $n>n_{0}$, we set $A_{n}^{(i)}$ to be the unique solution to the linear system (28) (with $n_{0}$ replaced by $n$) and $\psi_{n,n_{0}}^{*}A_{n}^{(i)}\varphi_{n,n_{0}}=A_{n_{0}}^{(i)}$. This linear system is typically large and sparse, and we solve it using LSQR [86]. The extension of $B_{n_{0}}^{(j)}$ is handled similarly, except that $n_{0}$ needs to exceed the presentation degrees of both $\mathscr{W}$ and $\mathscr{U}$ to guarantee that both $B_{n_{0}}^{(j)}$ and $(B_{n_{0}}^{(j)})^{*}$ extend to morphisms.
@@
We use the above algorithm to obtain dimension counts for parametric families of free descriptions. See the functions compute_dims_a, compute_dims_b, and compute_dims_c on GitHub for the code computing these dimensions using Algorithm 1.
@@
    The dimensions of invariants parametrizing free descriptions are $\dim\mathcal{L}(\mathbb{V}_{n},\mathbb{U}_{n})^{\mathsf{G}_{n}}=39$, $\dim\mathcal{L}(\mathbb{W}_{n},\mathbb{U}_{n})^{\mathsf{G}_{n}}\allowbreak=1068$, and $\dim\mathbb{U}_{n}^{\mathsf{G}_{n}}=17$ for $n\geq 8$. The dimensions of linear maps $\{A_{n}\}$ and $\{B_{n}\}$ satisfying Proposition 3.2(a), yielding intersection-compatible convex sets, are
@@
    $\dim\Big{\{}\{B_{n}\colon\mathbb{W}_{n}\to\mathbb{U}_{n}\}:\textrm{both }\{B_{n}\}\textrm{ and }\{B_{n}^{*}\}\textrm{ morphisms}\Big{\}}=104.$
@@
    $\dim\Big{\{}\{B_{n}\colon\mathbb{W}_{n}\to\mathbb{U}_{n}\}:\textrm{both }\{B_{n}\}\textrm{ and }\{B_{n}^{*}\}\textrm{ morphisms}\Big{\}}=37.$

### v3

@@ 4.7 Graphon Parameters
Figure 2: Weighted undirected graph represented as a graph, an adjacency matrix $X$, and a symmetric function (graphon) $W_{X}$ on $0,1]^{2}$.
@@
$\ +L_{4}\otimes\left(\mathrm{diag}(X)\mathbbm{1}^{\top}+\mathbbm{1}\mathrm{diag}(X)^{\top}\right)+L_{5}\otimes X+L_{6}\otimes\mathbbm{1}\mathbbm{1}^{\top}+L_{7}\otimes(2^{n}I_{2^{n}})\succeq 0\Bigg\},\qquad L_{7}\succeq 0.$
Note that all the functions of $X$ appearing in the above description can be interpreted as natural functions of the associated step graphon $W_{X}$. For example, $\frac{\mathbbm{1}^{\top}X\mathbbm{1}}{2^{2n}}=\int_{[0,1]^{2}}W_{X}(t,s)\,\mathrm{d}t\,\mathrm{d}s$ and $\frac{\mathrm{Tr}(X)}{2^{n}}=\int_{0}^{1}W_{X}(t,t)\,\mathrm{d}t$. Similarly, $\frac{1}{2^{n}}(X\mathbbm{1})_{i}=\int_{(i-1)/2^{n}}^{i/2^{n}}W_{X}((i-1/2)/2^{n},y)\,dy$ gives the values of the degree function of the graphon $W_{X}$, and $\mathrm{diag}(X)$ gives the values of the diagonal function $W_{X}(x,x)$ as $x$ ranges over $[0,1]$, see [15, Chap. 7].
Expressing the above dimension-free description in terms of graphons yields a description of a limiting convex set by Theorem 3.6. Its gauge function is, in turn, a convex graphon parameter extending continuously to the corresponding limit. Endow $\mathbb{V}_{\infty}$ with the $L^{\infty}([0,1]^{2})$-norm, noting that $\|\mathcal{P}_{n}W\|_{\infty}\leq\|W\|_{\infty}$ for all $W\in L^{\infty}([0,1]^{2})$ by Jensen’s inequality. We view elements $W_{i,j}(x,y)]_{i,j=1}^{k}\in\mathbb{S}^{k}\otimes L^{\infty}([0,1]^{2})$ as symmetric matrices with entries in $L^{\infty}([0,1]^{2})$ (so $W_{j,i}=W_{i,j}$), and denote
$W_{i,j}(x,y)]_{i,j=1}^{k}\succeq 0\quad\textrm{ if }\quad\int_{[0,1]^{2}}\sum_{i,j=1}^{k}W_{i,j}(x,y)f_{i}(x)f_{j}(y)\,dx\,dy\geq 0\quad\textrm{ for all }\quad f_{1},\ldots,f_{k}\in L^{2}([0,1]).$ (21)
Remarkably, convex graphon parameters that extend continuously to this limit are also projection-compatible by [77, Thm. 3.17].
@@
where $\overline{\mathbb{V}_{\infty}}\subseteq L^{\infty}([0,1]^{2})$ is a subspace of symmetric bounded measurable functions on $0,1]^{2}$.
@@
Endow $\mathbb{U}_{\infty}=\mathbb{S}^{k}\otimes\mathbb{V}_{\infty}$ with the norm $\|[W_{i,j}]_{i,j=1}^{k}\|=\max_{i,j}\|W_{i,j}\|_{\infty}$, so $\overline{\mathbb{U}_{\infty}}=\mathbb{S}^{k}\otimes\overline{\mathbb{V}_{\infty}}$. Then $\overline{\mathcal{K}_{\infty}}$ is the cone of PSD matrices $W_{i,j}(x,y)]_{i,j=1}^{k}\succeq 0$ in $\overline{\mathbb{U}_{\infty}}$. Furthermore, for any $W\in\mathbb{V}_{\infty}$ we have
@@ 5.1 The PSD Cone and Variants
$\mathbb{U}_{n}^{\mathsf{G}_{n}}=\left\{Q_{n}\left(\bigoplus_{\lambda\in\Lambda}(X_{\lambda}\otimes I_{\dim\mathbb{W}_{\lambda[n]}})\right)Q_{n}^{*}:X_{\lambda}\in\mathbb{S}^{m_{\lambda}}\right\},$ (24)
@@
$\mathrm{Sym}^{2}_{+}(\mathbb{V}_{n})^{\mathsf{G}_{n}}=\left\{X\in\mathbb{U}_{n}^{\mathsf{G}_{n}}:X\succeq 0\right\}=\left\{Q_{n}\left(\bigoplus_{\lambda\in\Lambda_{n}}(X_{\lambda}\otimes I_{\dim\mathbb{W}_{\lambda}})\right)Q_{n}^{*}:X_{\lambda}\in\mathbb{S}^{m_{\lambda}}_{+}\right\}.$ (25)
@@
We now prove Theorem 1.12 giving constant-sized descriptions for invariant sums-of-squares. s the vector whose coordinates are all the monomials in
@@
If $v(x)$ is the vector whose coordinates are all the monomials of degree at most $d$ in the variables $x_{i_{1},\ldots,i_{k}}$ with $1\leq i_{1}\leq\ldots\leq i_{k}\leq n$, then [82, Thm. 3.39] yields
$\mathrm{SOS}_{\mathbb{U}_{n}}=\pi_{n}\left(\mathrm{Sym}_{+}^{2}\left(\mathrm{Sym}^{\leq d}\left(\mathrm{Sym}^{k}\mathbb{R}^{n}\right)\right)\right),\quad\pi_{n}(M)=v(x)^{\top}Mv(x)+\mathcal{I}_{n}.$
The map $\pi_{n}\colon\mathrm{Sym}^{2}\left(\mathrm{Sym}^{\leq d}\left(\mathrm{Sym}^{k}\mathbb{R}^{n}\right)\right)\to\mathbb{U}_{n}$ is equivariant by definition of the action of $\mathsf{G}_{n}$ and the invariance of $\mathcal{I}_{n}$. If $\mathscr{V}_{0}=\{\mathbb{R}^{n}\}$ as in Example 1.5, then $\mathscr{V}=\mathrm{Sym}^{\leq d}\left(\mathrm{Sym}^{k}\mathscr{V}_{0}\right)$ is a $\mathscr{V}_{0}$-module generated in degree $kd$, and presented in degree $kd$ if $\mathsf{G}_{n}=\mathsf{S}_{n},\mathsf{B}_{n}$ and in degree $kd+1$ if $\mathsf{G}_{n}=\mathsf{D}_{n}$ by Theorem 2.11. Thus, the result follows from Theorem 5.1 and Proposition 5.3. ∎
@@ 5.2 The Relative Entropy Cone and Variants
Set $d_{0}=0$ if $\mathsf{G}_{n}=\mathsf{S}_{n},\mathsf{B}_{n}$ and $d_{0}=1$ if $\mathsf{G}_{n}=\mathsf{D}_{n}$. Let $\hat{\mathcal{B}}\subseteq\mathcal{B}_{d_{W}}$ be a set of minimal-degree $\mathsf{G}_{d_{W}+d_{0}}$-orbit representatives in $\mathcal{B}_{d_{W}+d_{0}}$, which are also orbit representatives for $\mathcal{B}_{n}$ for all $n\geq d_{W}+d_{0}$ by Proposition 2.15(c). Any $M\in\widetilde{\mathbb{U}_{n}}^{\mathsf{G}_{n}}=\mathcal{L}(\mathbb{R}^{\mathcal{B}_{n}},\mathbb{U}_{n})^{\mathsf{G}_{n}}$ for $n\geq d_{W}+d_{0}$ is fully determined by the images $Me_{\beta}\in\mathbb{U}_{n}^{\mathsf{Stab}_{\mathsf{G}_{n}}(\beta)}$ of the basis elements $e_{\beta}$ for $\beta\in\hat{\mathcal{B}}$ and conversely, for any collection $\left\{u_{\beta}\in\mathbb{U}_{n}^{\mathsf{Stab}_{\mathsf{G}_{n}}(\beta)}\right\}_{\beta\in\hat{\mathcal{B}}}$ there is a unique $M\in\widetilde{\mathbb{U}_{n}}^{\mathsf{G}_{n}}$ satisfying $Me_{\beta}=u_{\beta}$. Moreover, $M\in\widetilde{\mathcal{K}_{n}}$ if and only if $Me_{\beta}\in\mathcal{K}_{n}^{\mathsf{Stab}_{\mathsf{G}_{n}}(\beta)}$ for all $\beta\in\hat{\mathcal{B}}$. Thus, we have
@@
As an application of Proposition 5.5, we obtain constant-sized descriptions for SAGE cones of signomials. Indeed, if $\mathcal{A}_{n},\mathcal{B}_{n}\subseteq\mathbb{R}^{n}$ as in the above proposition, define the sequence $\mathscr{F}=\{\mathbb{F}_{n}\}$ of functions on $\{\mathbb{R}^{n}\}$
@@
$\mathrm{SAGE}_{\mathcal{A}_{n},\mathcal{B}_{n}}$ $=\Big\{(c,t)\in\mathbb{R}^{\mathcal{A}_{n}}\oplus\mathbb{R}^{\mathcal{B}_{n}}:\exists M=(\nu^{(\beta)},c^{(\beta)},t_{\beta})_{\beta\in\mathcal{B}_{n}}\in\mathrm{REM}_{\mathcal{A}_{n},\mathcal{B}_{n}}\textrm{ s.t. }\sum\nolimits_{\beta\in\mathcal{B}_{n}}c^{(\beta)}=c,$
$\hskip 18.49988pt\sum\nolimits_{\alpha\in\mathcal{A}_{n}}\nu^{(\beta)}_{\alpha}(\alpha-\beta)=0\textrm{ for all }\beta\in\mathcal{B}_{n}\Big\}$
@@
$\mathbb{L}_{n}$ $=\left\{M=(\nu^{(\beta)},c^{(\beta)},t_{\beta})_{\beta\in\mathcal{B}_{n}}\in\widetilde{\mathbb{U}_{n}}:\sum\nolimits_{\alpha\in\mathcal{A}_{n}}\nu^{(\beta)}_{\alpha}(\alpha-\beta)=0\textrm{ for all }\beta\in\mathcal{B}_{n}\right\},$
@@
6 Dimension-Free Convex Regression
In this section, we use our framework to describe a solution to the dimension-free convex regression problem. Recall that in this problem we are given evaluation data in different dimensions and we seek a sequence of convex functions that best fit the data and can be instantiated in any desired dimension (including those not represented in the data). To that end, we use the framework we developed in Section 3 to obtain parametric families of freely-described convex sets in a fully algorithmic manner. We then present a numerical procedure to fit elements of these families to the given data. Our implementation of the resulting algorithms can be found at https://github.com/eitangl/anyDimCvxSets.
@@ 6.1 Computationally Parametrizing Descriptions
1:Input: Finitely-presented consistent sequences $\mathscr{V},\mathscr{W},\mathscr{U}$.
@@
Fix $n_{0}\in\mathbb{N}$ as in Algorithm 1, computed for instance using Theorem 2.11. We now explain how to find bases for the desired spaces of equivariant linear maps $A_{n_{0}}$ and $B_{n_{0}}$ in a fixed dimension, and how to extend them to bases of freely-described elements of $\mathscr{V}\otimes\mathscr{U}$ and $\mathscr{W}\otimes\mathscr{U}$. These elaborate on steps 5, 8 and 10 in Algorithm 1.
@@
Let $\mathsf{G}$ be a real Lie group with finitely-many connected components acting on a vector space $\mathbb{V}$ via the homomorphism $\rho\colon\mathsf{G}\to\mathrm{GL}(\mathbb{V})$. Let $\{H_{i}\}$ be a basis for the Lie algebra $\mathfrak{g}$ of $\mathsf{G}$ and $\{h_{j}\}\subseteq\mathsf{G}$ be a finite collection such that $\mathsf{G}$ is generated by $\mathrm{exp}(\mathfrak{g})$ and $\{h_{j}\}$ [^6] Such $\{h_{j}\}$ exist by [31, Appdx. H].. Then
@@
The equations (28) express the space $\mathcal{L}(\mathbb{V}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ as the kernel of a matrix, which is often very large and sparse. Forming this matrix only requires specifying the matrices representing $\{H_{i}\}$ and $\{h_{j}\}$ for $\mathsf{G}_{n_{0}}$. Computing a basis for the kernel of this matrix can then by done using its LU decomposition as in [83, 84], or using the algorithm of [31, §5]. A basis for the space $\mathcal{L}(\mathbb{W}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ is obtained analogously.
@@
hence the corresponding subspace of $\mathcal{L}(\mathbb{V}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ is the kernel of the matrix obtained by combining (28)-(30). Forming this matrix requires specifying the matrices representing $\{H_{i}\}$ and $\{h_{j}\}$ for $\mathsf{G}_{n_{0}}$, as well as the matrices representing the embeddings $\varphi_{n_{0},i}$ and $\psi_{n_{0},i}$. The subspace of $\mathcal{L}(\mathbb{W}_{n_{0}},\mathbb{U}_{n_{0}})^{\mathsf{G}_{n_{0}}}$ satisfying the hypotheses of Theorem 3.5 is again the kernel of a matrix and its basis is computed similarly.
@@
Given bases of equivariant $A_{n_{0}}^{(i)},B_{n_{0}}^{(j)}$, we wish to extend them to freely-described elements. We do so computationally by applying a linear map for $n<n_{0}$ and solving a linear system for each $n>n_{0}$ to which we wish to extend. For each $n<n_{0}$, we set $A_{n}^{(i)}=\psi_{n_{0},n}^{*}A_{n_{0}}^{(i)}\varphi_{n_{0},n}$ and similarly for $B_{n}^{(j)}$. For each $n>n_{0}$, we set $A_{n}^{(i)}$ to be the unique solution to the linear system (28) (with $n_{0}$ replaced by $n$) and $\psi_{n,n_{0}}^{*}A_{n}^{(i)}\varphi_{n,n_{0}}=A_{n_{0}}^{(i)}$. Forming this linear system requires sepcifying the matrices representing $\{H_{i}\}$ and $\{h_{j}\}$ for $\mathsf{G}_{n}$, as well as matrices representing $\varphi_{n,n_{0}}$ and $\psi_{n,n_{0}}$. The above linear system is typically large and sparse, and we solve it using LSQR [85]. The extension of $B_{n_{0}}^{(j)}$ is handled similarly, except that $n_{0}$ needs to exceed the presentation degrees of both $\mathscr{W}$ and $\mathscr{U}$ to guarantee that both $B_{n_{0}}^{(j)}$ and $(B_{n_{0}}^{(j)})^{*}$ extend to morphisms.
@@
We use the above algorithm to obtain dimension counts for parametric families of dimension-free descriptions. See the functions compute_dims_a, compute_dims_b, and compute_dims_c on GitHub for the code computing these dimensions using Algorithm 1.
@@
    The dimensions of invariants parametrizing dimension-free descriptions are $\dim\mathcal{L}(\mathbb{V}_{n},\mathbb{U}_{n})^{\mathsf{G}_{n}}=39$, $\dim\mathcal{L}(\mathbb{W}_{n},\mathbb{U}_{n})^{\mathsf{G}_{n}}\allowbreak=1068$, and $\dim\mathbb{U}_{n}^{\mathsf{G}_{n}}=17$ for $n\geq 8$. The dimensions of linear maps $\{A_{n}\}$ and $\{B_{n}\}$ satisfying Proposition 3.2(a), yielding intersection-compatible convex sets, are
@@
[…]
@@
[…]